[figure: 9.7 cm]

“COMB-LIKE” POLARONS AND BIPOLARONS IN ANTIFERROMAGNETIC MATERIALS

E. Hankiewicz, R. Buczko, and Z. Wilamowski Institute of Physics, Polish Academy of Sciences, Warsaw, Poland

The concept of a magnetic bipolaron as a possible candidate for Cooper-like pair in HT_{c}- compounds is well known. Usually antiferromagnetic bipolarons were considered on the basis of a classical model of the antiferromagnet (for example Mott-like bipolarons). In contrast in the present study we indicate a possibility of a new type of antiferromagnetic bipolaron formation, in the quantum antiferromagnet (QAF) approach. We examine a Hamiltonian which is a sum of d-d interaction between localized d-spins in a finite 1D or 2D antiferromagnetic system and the sp-d interaction of QAF with a sp carrier. Varying the shape of spatial distribution of carrier we diagonalize the Hamiltonian strictly using the Lanczos method. The results show that the “comb-like” distribution of carrier density on every second spin of QAF is energetically the most favorable. In contrast to the Mott-like polaron which induces a local moment in the QAF system our new “comb-like” polaron orginates from damping of antiferromagnetic quantum fluctuations only. The most important feature of these polarons is a tendency to form pairs (bipolarons) glued by indirect exchange via QAF. We also present a possible scenario of the antiferromagnet - superconductor transition.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Ewelina Hankiewicz

Address for correspondence : Instytut Fizyki PAN al. Lotników 32/46 02–668 Warszawa

Email address : sabcia@ifpan.edu.pl